OpenStudy (anonymous):

Medal + Fan: The population of a town grows exponentially. After 1 year, the population is 34,560. After 2 years, the population is 37,325. Which equation can be used to predict, y, the number of people living in the town after x years? (Round population values to the nearest whole number.)

OpenStudy (anonymous):

y = 32,000(1.08)x y = 32,000(0.08)x y = 31,795(1.08)x y = 31,795(0.08)x

OpenStudy (anonymous):

Got to go with C

OpenStudy (wolf1728):

At year zero population = 32,000*(1.08)^0 Year one 32,000*(1.08)^1 Year two 32,000*(1.08)^2 So population = 32,000 * (1.08)^years I'd say the answer is A Also those answers are written very poorly answer A should read y = 32,000 * (1.08) ^ x